# QuSplit: Achieving Both High Fidelity and Throughput via Job Splitting on Noisy Quantum Computers

### v1

@@ Abstract
As we enter the quantum utility era, the computing paradigm shifts toward quantum-centric computing, where multiple quantum processors collaborate with classical computers, exemplified by platforms like IBM Quantum and Amazon Braket. In this paradigm, efficient resource management is crucial; however, unlike classical computing, quantum processors face significant challenges due to noise, which raises fidelity concerns in quantum applications. Compounding this issue, the noise characteristics across different quantum processors are inherently heterogeneous, making resource optimization even more complex. Existing resource management strategies primarily focus on mapping and scheduling jobs to these heterogeneous backends, which leads to some jobs suffering extremely low fidelity. Targeting quantum optimization jobs (e.g., VQC, VQE, QAOA) — one of the most promising quantum applications in the NISQ era, we hypothesize that running the later stages of a job on a high-fidelity quantum processor can significantly enhance overall fidelity. To validate this hypothesis, we use the VQE as a case study and propose a novel and efficient Genetic Algorithm-based scheduling framework with the consideration of job splitting. Experimental results demonstrate that our approach maintains high fidelity across all jobs and significantly improves system throughput. Furthermore, the proposed algorithm shows excellent scalability with respect to the number of quantum processors and the volume of jobs, making it a robust solution for emerging quantum computing platforms.
I Introduction
@@
Quantum computing holds great promise to revolutionize domain applications, such as computational chemistry [1, 2, 3, 4, 5, 6, 7, 8, 9] and material science [10, 11, 12, 13, 14, 15, 16, 17], and optimizations [18, 19, 20, 21, 22, 23, 24, 25, 26], outperforming state-of-the-art classical computing significantly in terms of speed, accuracy, or the ability to handle larger problem sizes. Along with the emergence of quantum computers, the computing paradigm gradually shifts toward quantum-centric computing, where multiple quantum processors collaborate with classical computers. These quantum processors are inherently heterogeneous due to the presence of different quantum bit (qubit) technologies, such as superconducting qubits (e.g., IBM, Regitte), trapped ion qubits (e.g., IonQ, Quantinuum) and neutral atoms (e.g., QuEra, Infleqtion), bringing quite varied noise levels. Even for platforms with the same qubit technology, say IBM Quantum, the noise levels are different from processor to processor. The emergence of quantum-centric computing and the presence of heterogeneity call for automated resource management with two fundamental needs: (1) high quantum processor utilization or job execution throughput and (2) high execution fidelity across quantum processors.
@@
As we are still in the infancy of quantum-centric computing, most vendors require users to specify quantum processors for their jobs, which manually significantly impedes domain scientists who lack quantum knowledge for quantum-augmented scientific discovery. Only until recently, there emerging works for automated parallel scheduling algorithm [27, 28]; however, these quantum job scheduling methods follow the traditional scheduling problem developed for classical computing, that is, they focus on assigning one entire job to a specific backend to satisfy the qubit constraints, minimizing makespans, and maximizing fidelity. Compared with the classical scheduling problem, the new additional objective on fidelity brings great challenges. Firstly, it creates the conflict to minimize makespans since every job desires the quantum processor with the lowest noise level for high fidelity. More importantly, secondly, it has a great pitfall if the optimizer uses average fidelity for optimization, which can easily lead to unfair treatment on different jobs, but a soundness average fidelity.
@@
II Background and Related Work
II-A Quantum-centric computing
Over the past decade, quantum computing has achieved tremendous advances in both hardware and software, including continuously increasing quantum bit (qubit) numbers and emerging applications from different domains. For the foreseeable future, a highly coupled and integrated infrastructure with quantum computers and classical computers will be necessary to ensure efficient and large-scale scientific research and innovation, known as Quantum-Centric Computing, which has emerged already, evidenced by IBM Quantum with 14 quantum processors [29] and Amazon Braket [30] with 7 quantum computers from IonQ, IQM, QuEra, and Rigetti.
II-B Quantum Jobs and Applications
@@
II-C Quantum Scheduling
In quantum-centric computing with multiple heterogeneous quantum processors, efficient scheduling methods for various quantum applications are essential to optimize the utilization of available quantum processors, especially as noise levels vary across different quantum devices. Several works have explored quantum scheduling strategies [31, 32, 33] aimed at optimizing backend assignments to improve fidelity or minimize execution time. Most recently, [27, 28] developed parallel scheduling algorithms. These methods typically focus on assigning entire quantum jobs to specific backends based on their characteristics, such as noise levels and qubit connectivity. Some other approaches also involve partitioning a single backend’s qubits to run multiple jobs concurrently for resource sharing, maximizing resource utilization [34, 35].
@@ III Observation and Motivation
III-A Observation 1: Different noise levels lead to different performance
TABLE I: Comparison of Backend Properties and VQE Results for the energy estimation of $H_{2}$ Molecular.
@@ III-A Observation 1: Different noise levels lead to different performance
Table I presents the VQE results for the $H_{2}$ molecular on two different backends: ibm_nazca and IonQ. Each backend exhibits unique noise characteristics, as reflected in their respective 1-qubit, 2-qubit, and readout errors on average in Table I. The IonQ backend, with significantly lower error rates (e.g., 5.10E-03 v.s. 3.12E-02 in readout error), achieves an approximate energy of -1.79, which is much closer to the reference value of -1.86, resulting in a smaller error of 0.07. In contrast, the ibm_nazca backend, with higher error rates, converges to an approximate energy of -1.37, leading to a larger error of 0.49. This demonstrates how backend noise impacts the fidelity of VQE results.
@@
Figure 1: Impact of job splitting on VQE energy and convergence efficiency.
Figure 2: Effect of tail stage length on VQE energy with job splitting method.
III-B Observation 2: Splitting jobs can help improve performance
@@
III-C Observation 3: Split ratio impacts the performance improvement
@@ III-D Motivation
IV QuSplit Framework
IV-A Design Philosophy of QuSplit
@@ IV-B Overview of QuSplit
Figure 3: Overview of the proposed QuSplit framework.
@@
$score(B)=\frac{\sum_{j\in\mathcal{J}}score(j,B)}{|\mathcal{J}|}$ (2)
@@
IV-C Fidelity-aware Quantum Scheduler
@@
$Sol^{*}=\arg\max_{\forall Sol}\left(w_{1}\cdot TH(Sol)+w_{2}\cdot FI(Sol)\right)$ (3)
@@
V Experimental Results
TABLE II: Performance Scores of VQE Jobs with Split and Non-Split Configurations Across Different Molecules.
@@ V-A Experimental Setup
Figure 4: Design space exploration of job scheduling strategies for VQE jobs using QuSplit.
Figure 5: QuSplit evaluation on large-scale job scheduling scenarios with the split ratio of 0.2, 0.4, 0.6, 0.8 compared with all executions on the best backend (Method 1) and backend round-robin scheduling (Method 2).
@@
V-B Evaluation of Job-Splitting
Table II summarizes the performance of VQE jobs for three different molecules ($H_{2}$, $BF$, and $CO_{2}$) under various scheduling strategies, including non-split and split configurations (SplitB1_B3) with different split ratios ($r$ = 0.4 and $r$ = 0.6). Backend B3 is better than B1. The results demonstrate the following key observations:
@@ V-B Evaluation of Job-Splitting
Performance Improvement through Job-Splitting. Job-splitting significantly enhances performance compared to running the entire job on the noisier backend (B1). For example, for $H_{2}$, the performance improves from 0.79 (B1) to 0.98 under both split configurations ($r$ = 0.4 and $r$ = 0.6). Similar improvements are seen across the other molecules.
@@
Effect of Split Ratio. Allocating a larger proportion of iterations to the higher-fidelity backend (B3) usually yields better performance. For instance, the $BF$ molecule achieves a performance score of 0.78 with a split ratio of $r$ = 0.4, which increases to 0.86 with $r$ = 0.6. This trend indicates that the split ratio plays a crucial role in balancing performance and resource utilization.
@@
Comparison with All B3 Strategy. While executing all iterations on B3 generally yields the best performance (e.g., 0.95 for $BF$ and 0.94 for $CO_{2}$), the split strategy offers a practical alternative. By leveraging the lower-fidelity backend (B1) in the early stages, job-splitting achieves comparable or near-optimal performance.
V-C Design Space Exploration
@@
Figure 4 showcases the design space exploration for scheduling strategies involving 2 VQE jobs (Figure 4(a)) and 3 VQE jobs (Figure 4(b)) across three backends, representing small-scale job scheduling scenarios. The performance score (y-axis) and throughput (x-axis) trade-off for all possible strategies are represented as blue dots, with the strategies identified by QuSplit highlighted in red circles.
@@
For 2 VQE jobs, brute-force enumeration of all possible strategies is computationally feasible, allowing the entire design space to be explored. QuSplit effectively identifies high-quality strategies (red circles) that lie on or near the Pareto Frontier, demonstrating its capability to select optimal trade-offs between performance and throughput. Similarly, for 3 VQE jobs, although the design space grows, brute-force exploration is still manageable, while QuSplit continues to identify optimal strategies.
V-D QuSplit Evaluation on Large-Scale Job Scenarios
@@
Figure 5 illustrates the relationship between throughput and average performance score for various job scheduling strategies, evaluated across different job sizes ranging from 10 to 80 jobs, representing large-scale job scheduling scenarios. The strategies include:
@@
  • Method 1 (red marker): All jobs are executed on the highest-fidelity backend (B3), prioritizing performance over throughput.
  • Method 2 (purple marker): Round-robin scheduling [36, 37], which is widely studied in classical and cloud computers, across all backends in the sequence of $\langle B3,B2,B1,\cdots\rangle$.
@@
  • Method 1 achieves the highest performance but suffers from low throughput, making it less practical for large workloads.
  • Method 2 delivers better throughput but at the expense of significantly lower performance.
  • As the number of jobs increases, QuSplit strategies maintain their advantage, demonstrating scalability even with 80 jobs.
V-E Scheduling Time Analysis
Figure 6: Elapsed time of QuSplit for different numbers of jobs across 3 and 10 backends.
@@ V-E Scheduling Time Analysis
The scheduling time for 10 backends (black line) is consistently higher than for 3 backends (blue line). This is due to the increased complexity of searching for optimal job allocations as the number of available backends grows. The additional backends expand the solution space, leading to longer computation times. Both configurations show a linear increase in scheduling time as the number of jobs grows from 10 to 80. However, even with 10 backends, the scheduling time remains within acceptable limits, demonstrating the scalability of the scheduling algorithm.
V-F Validation of Backend Fidelity Score
Figure 7: Validating benchmarks-based backends fidelity score is proportional to the fidelity of quantum VQE jobs.
Figure 7 shows the fidelity scores of five random benchmark circuits and the performance scores of three VQE jobs evaluated on three backends: B1, B2, and B3. Experimental results indicate that backend B3 consistently achieves the highest scores across all circuits, followed by B2 and B1, demonstrating an intuition of lower noise levels correlating with higher performance. For example, Job 1 achieves a score of 0.911 on B3, significantly outperforming its scores on B2 (0.732) and B1 (0.679). Similar patterns are observed for Job 2 and Job 3.

### v2

@@ Abstract.
With the progression into the quantum utility era, computing is shifting toward quantum-centric architectures, where multiple quantum processors collaborate with classical computing resources. Platforms such as IBM Quantum and Amazon Braket exemplify this trend, enabling access to diverse quantum backends. However, efficient resource management remains a challenge, as quantum processors are highly susceptible to noise, which significantly impacts computation fidelity. Additionally, the heterogeneous noise characteristics across different processors add further complexity to scheduling and resource allocation. Existing scheduling strategies typically focus on mapping and scheduling jobs to these heterogeneous backends, which leads to some jobs suffering extremely low fidelity. Targeting quantum optimization jobs (e.g., VQC, VQE, QAOA) — among the most promising quantum applications in the NISQ era — we hypothesize that executing the later stages of a job on a high-fidelity quantum processor can significantly improve overall fidelity. To verify this, we use VQE as a case study and develop a Genetic Algorithm-based scheduling framework that incorporates job splitting to optimize fidelity and throughput. Experimental results demonstrate that our approach consistently maintains high fidelity across all jobs while significantly enhancing system throughput. Furthermore, the proposed algorithm exhibits excellent scalability in handling an increasing number of quantum processors and larger workloads, making it a robust and practical solution for emerging quantum computing platforms. To further substantiate its effectiveness, we conduct experiments on a real quantum processor, IBM Strasbourg, which confirm that job splitting improves fidelity and reduces the number of iterations required for convergence.
copyright: nonecopyright: none
1. Introduction
@@
Quantum computing holds great promise to revolutionize domain applications, such as computational chemistry (Motta and Rice, 2022; Von Lilienfeld, 2018; Levine et al., 2009; Shao et al., 2006; Feniou et al., 2023; Cao et al., 2019; Peruzzo et al., 2014; Izmaylov et al., 2019; Izsák et al., 2023) and material science (Bauer et al., 2020; Alexeev et al., 2024; Teale et al., 2022; Bochevarov et al., 2013; Hafner, 2007; Ebbesen, 2016; Reiser et al., 2022; Westermayr et al., 2021), and optimizations (Jiang et al., 2021a; Wang et al., 2021; Jiang et al., 2021b; Stein et al., 2022; Jing et al., 2022; Zeng et al., 2022; Tacchino et al., 2020; Yan et al., 2020; Gratsea and Huembeli, 2022), outperforming state-of-the-art classical computing significantly in terms of speed, accuracy, or the ability to handle larger problem sizes. Along with the emergence of quantum computers, the computing paradigm gradually shifts toward quantum-centric computing, where multiple quantum processors collaborate with classical computers. These quantum processors are inherently heterogeneous due to the presence of different quantum bit (qubit) technologies, such as superconducting qubits (e.g., IBM, Regitte), trapped ion qubits (e.g., IonQ, Quantinuum) and neutral atoms (e.g., QuEra, Infleqtion), bringing quite varied noise levels. Even for platforms with the same qubit technology, say IBM Quantum, the noise levels are different from processor to processor. The emergence of quantum-centric computing and the presence of heterogeneity call for automated resource management with two fundamental needs: (1) high quantum processor utilization or job execution throughput and (2) high execution fidelity across quantum processors.
@@
As we are still in the infancy of quantum-centric computing, most vendors require users to specify quantum processors for their jobs, which manually significantly impedes domain scientists who lack quantum knowledge for quantum-augmented scientific discovery. Only until recently, there emerging works for automated parallel scheduling algorithm (Orenstein and Chaudhary, 2024; Liu and Dou, 2024); however, these quantum job scheduling methods follow the traditional scheduling problem developed for classical computing, that is, they focus on assigning one entire job to a specific backend to satisfy the qubit constraints, minimizing makespans, and maximizing fidelity. Compared with the classical scheduling problem, the new additional objective on fidelity brings great challenges. Firstly, it creates the conflict to minimize makespans since every job desires the quantum processor with the lowest noise level for high fidelity. More importantly, secondly, it has a great pitfall if the optimizer uses average fidelity for optimization, which can easily lead to unfair treatment on different jobs, but a soundness average fidelity.
@@
2. Background and Related Work
2.1. Quantum-centric computing
Over the past decade, quantum computing has achieved tremendous advances in both hardware and software, including continuously increasing quantum bit (qubit) numbers and emerging applications from different domains. For the foreseeable future, a highly coupled and integrated infrastructure with quantum computers and classical computers will be necessary to ensure efficient and large-scale scientific research and innovation, known as Quantum-Centric Computing, which has emerged already, evidenced by IBM Quantum with 14 quantum processors (IBM, 2024) and Amazon Braket (Services, 2024) with 7 quantum computers from IonQ, IQM, QuEra, and Rigetti.
2.2. Quantum Jobs and Applications
@@
2.3. Quantum Scheduling
In quantum-centric computing with multiple heterogeneous quantum processors, efficient scheduling methods for various quantum applications are essential to optimize the utilization of available quantum processors, especially as noise levels vary across different quantum devices. Several works have explored quantum scheduling strategies (Ravi et al., 2021; Romero-Alvarez et al., 2024; Zhang et al., 2022) aimed at optimizing backend assignments to improve fidelity or minimize execution time. Most recently, (Orenstein and Chaudhary, 2024; Liu and Dou, 2024) developed parallel scheduling algorithms. These methods typically focus on assigning entire quantum jobs to specific backends based on their characteristics, such as noise levels and qubit connectivity. Some other approaches also involve partitioning a single backend’s qubits to run multiple jobs concurrently for resource sharing, maximizing resource utilization (Liu and Dou, 2021; Niu and Todri-Sanial, 2023).
@@ 3. Observation and Motivation
3.1. Observation 1: Different noise levels lead to different performance
Table 1. Comparison of Backend Properties and VQE Results for the energy estimation of $H_{2}$ Molecular.
@@ 3.1. Observation 1: Different noise levels lead to different performance
Table 1 presents the VQE results for the $H_{2}$ molecular on two different backends: ibm_nazca and IonQ. Each backend exhibits unique noise characteristics, as reflected in their respective 1-qubit, 2-qubit, and readout errors on average in Table 1. The IonQ backend, with significantly lower error rates (e.g., 5.10E-03 v.s. 3.12E-02 in readout error), achieves an approximate energy of -1.79, which is much closer to the reference value of -1.86, resulting in a smaller error of 0.07. In contrast, the ibm_nazca backend, with higher error rates, converges to an approximate energy of -1.37, leading to a larger error of 0.49. This demonstrates how backend noise impacts the fidelity of VQE results.
@@
Figure 1. Impact of job splitting on VQE energy and convergence efficiency.
Figure 2. Effect of tail stage length on VQE energy with job splitting method.
3.2. Observation 2: Splitting jobs can help improve performance
@@
3.3. Observation 3: Split ratio impacts the performance improvement
@@ 3.4. Motivation
4. QuSplit Framework
4.1. Design Philosophy of QuSplit
@@ 4.2. Overview of QuSplit
Figure 3. Overview of the proposed QuSplit framework.
@@
(2) $score(B)=\frac{\sum_{j\in\mathcal{J}}score(j,B)}{|\mathcal{J}|}$
@@
4.3. Fidelity-aware Quantum Scheduler
@@
(3) $Sol^{*}=\arg\max_{\forall Sol}\left(w_{1}\cdot TH(Sol)+w_{2}\cdot FI(Sol)\right)$
@@
4.4. Example of Scheduling Strategy
Figure 4. Example of job assignments for a scheduling strategy.
Figure 4 illustrates how quantum jobs are assigned to different backends under a specific scheduling strategy for 5 jobs with 3 available backends. The strategy used in this example is as follows: [$split_{B2B3}$, $split_{B1B2}$, $B3$, $split_{B2B3}$, $split_{B1B2}$] This strategy specifies how each job is executed, including whether it is split between two backends and which backends are used for each stage.
Example Job Input: The top section of the figure shows five quantum jobs (Job 1 to Job 5) and a predefined split ratio of 0.5. This split ratio indicates that for split jobs, 50% of the iterations will run on the first backend (Stage 1) and the remaining 50% will run on the second backend (Stage 2).
Job Assignments: The bottom section of the figure illustrates how the stages of each job are mapped to the available backends (Backend 1, Backend 2, and Backend 3) according to the specified strategy:
  • Job 1 ($split_{B2B3}$): Stage 1 is assigned to Backend 2, and Stage 2 is assigned to Backend 3.
  • Job 2 ($split_{B1B2}$): Stage 1 is assigned to Backend 1, and Stage 2 to Backend 2.
  • Job 3 ($B3$): The entire job runs on Backend 3.
  • Job 4 ($split_{B2B3}$): Stage 1 is assigned to Backend 2, and Stage 2 to Backend 3.
  • Job 5 ($split_{B1B2}$): Stage 1 is assigned to Backend 1, and Stage 2 to Backend 2.
This example demonstrates the flexibility of QuSplit in distributing job stages across backends to balance performance and resource utilization. The split strategy allows leveraging high-fidelity backends for critical computations while utilizing lower-fidelity backends for non-critical stages.
5. Experimental Results
Table 2. Performance Scores of VQE Jobs with Split and Non-Split Configurations Across Different Molecules.
@@ 5.1. Experimental Setup
Figure 5. Design space exploration of job scheduling strategies for VQE jobs using QuSplit.
Figure 6. QuSplit evaluation on large-scale job scheduling scenarios with the split ratio of 0.2, 0.4, 0.6, 0.8 compared with all executions on the best backend (Method 1) and backend round-robin scheduling (Method 2).
@@
5.2. Evaluation of Job-Splitting
Table 2 summarizes the performance of VQE jobs for three different molecules ($H_{2}$, $BF$, and $CO_{2}$) under multiple scheduling strategies. These include both non-split executions, where the entire job is run on a single backend, and job-splitting configurations (SplitB1_B3) with different split ratios ($r$ = 0.4 and $r$ = 0.6). Here, B3 represents a higher-fidelity quantum processor, whereas B1 is more error-prone. The results illustrate how job splitting influences performance under varying execution conditions, leading to the following key observations.
@@ 5.2. Evaluation of Job-Splitting
Performance Improvement through Job-Splitting. Job splitting proves to be a highly effective technique for improving VQE performance compared to executing the entire computation on the noisier backend (B1). For example, in the case of $H_{2}$, the performance improves from 0.79 (B1) to 0.98 under both split configurations ($r$ = 0.4 and $r$ = 0.6). Similar improvements are seen across the other molecules, suggesting that incorporating a more reliable backend in the latter stages of execution helps mitigate the adverse effects of noise and improve performance.
@@
Effect of Split Ratio. The proportion of iterations allocated to the high-fidelity backend (B3) plays a crucial role in determining the final performance. A larger allocation to B3 generally yields better performance. For instance, the VQE performance for the $BF$ molecule achieves a performance score of 0.78 with a split ratio of $r$ = 0.4, which increases to 0.86 with $r$ = 0.6. This pattern reinforces the importance of optimizing the split ratio to strike a balance between leveraging computational efficiency and maximizing result accuracy.
@@
Comparison with All B3 Strategy. While executing all iterations on B3 generally yields the best performance (e.g., 0.95 for $BF$ and 0.94 for $CO_{2}$), the split strategy offers a practical alternative. By leveraging the lower-fidelity backend (B1) in the early stages, job-splitting achieves comparable or near-optimal performance. This indicates that job-splitting can be a practical scheduling strategy, particularly when access to high-fidelity quantum processors is constrained.
5.3. Design Space Exploration
@@
Figure 5 showcases the design space exploration for scheduling strategies involving two VQE jobs (Figure 5(a)) and three VQE jobs (Figure 5(b)) across three quantum backends. These experiments represent small-scale job scheduling scenarios, where various execution strategies are evaluated based on their performance score (y-axis) and throughput (x-axis). Each blue dot in the figure corresponds to a unique scheduling strategy, while the strategies identified by QuSplit are highlighted using red circles.
@@
For two VQE jobs, brute-force enumeration of all possible strategies is computationally feasible, allowing the entire design space to be explored. QuSplit effectively identifies high-quality strategies (red circles) that lie on or near the Pareto Frontier, demonstrating its ability to identify optimal trade-offs between performance and throughput. Similarly, when scaling up to three VQE jobs, the design space expands, increasing the number of possible scheduling strategies. Despite this growth, brute-force exploration is still manageable, allowing a full assessment of the landscape. Notably, QuSplit continues to identify near-optimal scheduling configurations, reinforcing its effectiveness in balancing performance and execution efficiency even as the scheduling complexity increases.
5.4. QuSplit Evaluation on Large-Scale Job Scenarios
@@
Figure 6 illustrates the relationship between throughput and average performance score for various job scheduling strategies, evaluated across different job sizes ranging from 10 to 80 jobs. These experiments represent large-scale job scheduling scenarios, where different methodologies are compared in terms of their ability to balance computational efficiency and execution fidelity. The strategies analyzed include:
@@
  • Method 1 (red marker): All jobs are executed exclusively on the highest-fidelity backend (B3), maximizing performance at the cost of reduced throughput.
  • Method 2 (purple marker): Round-robin scheduling (Rasmussen and Trick, 2008; Balharith and Alhaidari, 2019), which is widely studied in classical and cloud computing, where jobs are cyclically assigned to backends in the sequence of $\langle B3,B2,B1,\dots\rangle$.
@@
  • Method 1 consistently delivers the highest performance, as expected, but suffers from severely limited throughput, making it less practical for handling large workloads efficiently.
  • Method 2 improves throughput by distributing jobs more evenly, yet this comes at the cost of significantly lower performance due to its fixed allocation across backends with varying noise levels.
  • As the number of jobs increases, QuSplit strategies continue to scale effectively, retaining their advantage in balancing performance and execution efficiency even at 80 jobs.
These results highlight the practicality of QuSplit as a scheduling approach for large-scale quantum workloads. By intelligently distributing job execution across backends based on noise-aware split ratios, QuSplit enables efficient resource utilization while preserving computational accuracy.
5.5. Scheduling Time Analysis
Figure 7. Elapsed time of QuSplit for different numbers of jobs across 3 and 10 backends.
@@ 5.5. Scheduling Time Analysis
The results indicate that the scheduling time for 10 backends (black line) is consistently higher than for 3 backends (blue line). This increase is attributed to the growing complexity of searching for optimal job allocations as the number of available backends expands. With more backends, the solution space enlarges, requiring additional computation time to evaluate and determine efficient scheduling strategies. Despite this complexity, both configurations exhibit a linear increase in scheduling time as the number of jobs grows from 10 to 80. However, even with 10 backends, the scheduling time remains within acceptable limits, demonstrating the scalability of the scheduling algorithm.
5.6. Validation of Backend Fidelity Score
Figure 8. Validating benchmarks-based backends fidelity score is proportional to the fidelity of quantum VQE jobs.
Figure 8 shows the fidelity scores of five random benchmark circuits and the performance scores of three VQE jobs evaluated on three backends: B1, B2, and B3. Experimental results indicate that backend B3 consistently achieves the highest scores across all circuits, followed by B2 and B1, demonstrating an intuition of lower noise levels correlating with higher performance. Score(B) metric, shown in Figure 8, represents the proxy backend fidelity score computed according to Eq. 2. This score serves as a numerical estimation of each backend’s expected performance, derived from benchmark circuit results.
For example, Job 1 achieves a score of 0.911 on B3, significantly outperforming its scores on B2 (0.732) and B1 (0.679). Similar patterns are observed for Job 2 and Job 3.